\begin{thebibliography}{99}
\bibitem{Froe94}
   J.\ Froehlich,
   Parameter derivatives of the Jacobi polynomials and the Gaussian
   hypergeometric function,
   Integral Transforms Spec.\ Funct.\ 2 (1994) 253
\bibitem{Koep97}
   W.\ Koepf,
   Identities for families of orthogonal polynomials and special
   functions,
   Integral Transforms Spec.\ Funct.\ 5 (1997) 69
\bibitem{Szmy09a}
   R.\ Szmytkowski,
   A note on parameter derivatives of classical orthogonal
   polynomials,
   arXiv:0901.2639
\bibitem{Akta14}
   R.\ Akta{\c{s}},
   A note on parameter derivatives of the Jacobi polynomials on the
   triangle,
   Appl.\ Math.\ Comput.\ 247 (2014) 368
\bibitem{Akta15}
   R.\ Akta{\c{s}},
   On parameter derivatives of a family of polynomials in two
   variables,
   Appl.\ Math.\ Comput.\ 256 (2015) 769
\bibitem{Akta16}
   R.\ Akta{\c{s}},
   Representations for parameter derivatives of some Koornwinder
   polynomials in two variables,
   J.\ Egyptian Math.\ Soc.\ 24 (2016) 555
\bibitem{Bryc05}
   Yu.\ A.\ Brychkov, K.\ O.\ Geddes,
   On the derivatives of the Bessel and Struve functions with respect
   to the order,
   Integral Transforms Spec.\ Funct.\ 16 (2005) 187
\bibitem{Sesm14}
   J.\ Sesma,
   Derivatives with respect to the order of the Bessel function of the
   first kind,
   arXiv:1401.4850
\bibitem{Duns15}
   T.\ M.\ Dunster,
   On the order derivatives of Bessel functions,
   Constr.\ Approx.\ (2016) doi:10.1007/s00365-016-9355-1
\bibitem{Bryc16}
   Yu.\ A.\ Brychkov,
   Higher derivatives of the Bessel functions with respect to the 
   order,
   Integral Transforms Spec.\ Funct.\ 27 (2016) 566
\bibitem{Szmy06}
   R.\ Szmytkowski,
   On the derivative of the Legendre function of the first kind with
   respect to its degree,
   J.\ Phys.\ A 39 (2006) 15147
   [corrigendum: J.\ Phys.\ A 40 (2007) 7819]
\bibitem{Szmy07}
   R.\ Szmytkowski,
   Addendum to `On the derivative of the Legendre function of the 
   first kind with respect to its degree',
   J.\ Phys.\ A 40 (2007) 14887
\bibitem{Szmy09b}
   R.\ Szmytkowski,
   On the derivative of the associated Legendre function of the first
   kind of integer degree with respect to its order (with
   applications to the construction of the associated Legendre
   function of the second kind of integer degree and order),
   J.\ Math.\ Chem.\ 46 (2009) 231
\bibitem{Bryc10}
   Yu.\ A.\ Brychkov, 
   On the derivatives of the Legendre functions $P_{\nu}^{\mu}(z)$
   and $Q_{\nu}^{\mu}(z)$ with respect to $\mu$ and $\nu$, 
   Integral Transforms Spec.\ Funct.\ 21 (2010) 175 
\bibitem{Cohl10}
   H.\ S.\ Cohl,
   Derivatives with respect to the degree and order of associated
   Legendre functions for $|z|>1$ using modified Bessel functions,
   Integral Transforms Spec.\ Funct.\ 21 (2010) 581
\bibitem{Cohl11}
   H.\ S.\ Cohl,
   On parameter differentiation for integral representations of
   associated Legendre functions,
   Sym.\ Integ.\ Geom.:\ Meth.\ Appl.\ (SIGMA) 7 (2011) 050
\bibitem{Szmy11}
   R.\ Szmytkowski,
   On the derivative of the associated Legendre function of the first
   kind of integer order with respect to its degree (with
   applications to the construction of the associated Legendre
   function of the second kind of integer degree and order), 
   J.\ Math.\ Chem.\ 49 (2011) 1436
\bibitem{Szmy12}
   R.\ Szmytkowski,
   On parameter derivatives of the associated Legendre function of
   the first kind (with applications to the construction of the
   associated Legendre function of the second kind of integer degree
   and order),
   J.\ Math.\ Anal.\ Appl.\ 386 (2012) 332
\bibitem{Abad03}
   J.\ Abad, J.\ Sesma,
   Successive derivatives of Whittaker functions with respect to the
   first parameter,
   Comput.\ Phys.\ Commun.\ 156 (2003) 13
\bibitem{Anca08}
   L.\ U.\ Ancarani, G.\ Gasaneo,
   Derivatives of any order of the confluent hypergeometric function
   ${}_{1}F_{1}(a,b,z)$ with respect to the parameter $a$ or $b$,
   J.\ Math.\ Phys.\ 49 (2008) 063508
\bibitem{Anca09}
   L.\ U.\ Ancarani, G.\ Gasaneo,
   Derivatives of any order of the Gaussian hypergeometric function
   ${}_{2}F_{1}(a,b,c;z)$ with respect to the parameters $a$, $b$ and
   $c$,
   J.\ Phys.\ A 42 (2009) 395208
\bibitem{Anca10}
   L.\ U.\ Ancarani, G.\ Gasaneo,
   Derivatives of any order of the hypergeometric function
   ${}_{p}F_{q}(a_{1},\ldots,a_{p};b_{1},\ldots,b_{q};z)$ with respect
   to the parameters $a_{i}$ and $b_{i}$,
   J.\ Phys.\ A 43 (2010) 085210
\bibitem{Lewi81}
   L.\ Lewin,
   Polylogarithms and Associated Functions,
   North-Holland, New York, 1981
\bibitem{Apos10}
   T.\ M.\ Apostol,
   Zeta and related functions,
   in:\ NIST Handbook of Mathematical Functions,
   edited by F.\ W.\ J.\ Olver, D.\ W.\ Lozier, R.\ F.\ Boisvert and
   C.\ W.\ Clark,
   Cambridge University Press, Cambridge, 2010
   (Sec.\ 25.12)
\bibitem{Szmy13}
   R.\ Szmytkowski,
   The derivatives $[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=0}$ 
   and $[\partial^{3}P_{\nu}(z)/\partial\nu^{3}]_{\nu=0}$, where
   $P_{\nu}(z)$ is the Legendre function of the first kind,
   arXiv:1301.6586
\bibitem{Laur15}
   B.\ J.\ Laurenzi,
   Derivatives with respect to the order of the Legendre polynomials,
   arXiv:1502.06507
\bibitem{Magn66}
   W.\ Magnus, F.\ Oberhettinger, R.\ P.\ Soni,
   Formulas and Theorems for the Special Functions of Mathematical
   Physics, 3rd ed.,
   Springer, Berlin, 1966
\end{thebibliography}

\documentclass[10pt,draft]{article}
\usepackage{amsfonts,amsmath,amssymb,cite}
\newcommand{\Imag}{\mathop{\textrm{Im}}\nolimits}
\newcommand{\Li}{\mathop{\textrm{Li}}\nolimits}
\renewcommand{\theequation}{\thesection.\arabic{equation}}
\begin{document}
\title{On the derivatives
$\partial^{2}P_{\nu}(z)/\partial\nu^{2}$ and $\partial
Q_{\nu}(z)/\partial\nu$\\ 
of the Legendre functions with respect to their degrees}
\author{Rados{\l}aw Szmytkowski \\*[3ex]
Atomic and Optical Physics Division, \\
Department of Atomic, Molecular and Optical Physics, \\
Faculty of Applied Physics and Mathematics,
Gda{\'n}sk University of Technology, \\
ul.\ Gabriela Narutowicza 11/12, 
80--233 Gda{\'n}sk, Poland \\
email: radoslaw.szmytkowski@pg.edu.pl}
\date{}
\maketitle
\begin{center}
\textbf{Published as: Integral Transforms Spec.\ Funct. 28 (2017)
645--62} \\*[1ex]
\textbf{doi: 10.1080/10652469.2017.1339039} \\*[5ex]
\end{center}
\begin{abstract} 
We provide closed-form expressions for the degree-derivatives
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$ and $[\partial
Q_{\nu}(z)/\partial\nu]_{\nu=n}$, with $z\in\mathbb{C}$ and
$n\in\mathbb{N}_{0}$, where $P_{\nu}(z)$ and $Q_{\nu}(z)$ are the
Legendre functions of the first and the second kind, respectively.
For $[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$, we find that
\begin{displaymath}
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=n}
=-2P_{n}(z)\Li_{2}\frac{1-z}{2}+B_{n}(z)\ln\frac{z+1}{2}+C_{n}(z),
\end{displaymath}
where $\Li_{2}[(1-z)/2]$ is the dilogarithm function, $P_{n}(z)$ is
the Legendre polynomial, while $B_{n}(z)$ and $C_{n}(z)$ are certain
polynomials in $z$ of degree $n$. For $[\partial
Q_{\nu}(z)/\partial\nu]_{\nu=n}$ and $z\in\mathbb{C}\setminus[-1,1]$,
we derive 
\begin{eqnarray*}
\frac{\partial Q_{\nu}(z)}{\partial\nu}\bigg|_{\nu=n} 
&=& -P_{n}(z)\Li_{2}\frac{1-z}{2}
-\frac{1}{2}P_{n}(z)\ln\frac{z+1}{2}\ln\frac{z-1}{2}
+\frac{1}{4}B_{n}(z)\ln\frac{z+1}{2} 
\nonumber \\
&& -\,\frac{(-1)^{n}}{4}B_{n}(-z)\ln\frac{z-1}{2}
-\frac{\pi^{2}}{6}P_{n}(z)
+\frac{1}{4}C_{n}(z)-\frac{(-1)^{n}}{4}C_{n}(-z).
\end{eqnarray*}
A counterpart expression for $[\partial
Q_{\nu}(x)/\partial\nu]_{\nu=n}$, applicable when $x\in(-1,1)$, is
also presented. Explicit representations of the polynomials
$B_{n}(z)$ and $C_{n}(z)$ as linear combinations of the Legendre
polynomials are given.
\vskip3ex
\noindent
\textbf{Key words:} Legendre functions; parameter derivatives;
dilogarithm
\vskip1ex
\noindent
\textbf{MSC2010:} 33C05, 33B30
\end{abstract}
%
%
\section{Introduction}
\label{I}
\setcounter{equation}{0}
Over the past 20 years or so, a growth of interest in parameter
derivatives of various special functions has been observed. The
research done on the subject is documented in a number of papers
reporting diverse methods for finding such derivatives for orthogonal
polynomials in one \cite{Froe94,Koep97,Szmy09a} and two
\cite{Akta14,Akta15,Akta16} variables, for Bessel functions
\cite{Bryc05,Sesm14,Duns15,Bryc16}, for Legendre and allied functions
\cite{Szmy06,Szmy07,Szmy09b,Bryc10,Cohl10,Cohl11,Szmy11,Szmy12}, and
also for various types of hypergeometric functions
\cite{Froe94,Abad03,Anca08,Anca09,Anca10}.

In Refs.\ \cite{Szmy06,Szmy07}, we presented results of our
investigations on the first-order derivative of the Legendre function
of the first kind with respect to its degree. We showed that
$[\partial P_{\nu}(z)/\partial\nu]_{\nu=n}$, with
$n\in\mathbb{N}_{0}$, is of the form
\begin{equation}
\frac{\partial P_{\nu}(z)}{\partial\nu}\bigg|_{\nu=n}
=P_{n}(z)\ln\frac{z+1}{2}+R_{n}(z),
\label{1.1}
\end{equation}
where $P_{n}(z)$ is the Legendre polynomial of degree $n$ and
$R_{n}(z)$ is another polynomial in $z$ of the same degree. We
investigated properties of the polynomials $R_{n}(z)$ and arrived at
their several explicit representations, including the following one:
\begin{equation} 
R_{n}(z)=2[\psi(2n+1)-\psi(n+1)]P_{n}(z)
+2\sum_{k=0}^{n-1}(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}P_{k}(z),
\label{1.2} 
\end{equation}
where $\psi(z)=\mathrm{d}\ln\Gamma(z)/\mathrm{d}z$ is the digamma
function.

In the year 2012, Dr.\ George P.\ Schramkowski kindly informed the
present author that in the course of doing research on a certain
problem in theoretical hydrodynamics, he had come across higher-order
derivatives $[\partial^{k}P_{\nu}(z)/\partial\nu^{k}]_{\nu=n}$, with
$n\in\mathbb{N}_{0}$ and $k\geqslant2$. Using \emph{Mathematica\/},
Schramkowski found that
\begin{equation}
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=0}
=-2\Li_{2}\frac{1-z}{2},
\label{1.3}
\end{equation}
where 
\begin{equation}
\Li_{2}z=-\int_{0}^{z}\mathrm{d}t\:\frac{\ln(1-t)}{t}
\label{1.4}
\end{equation}
is the dilogarithm function \cite{Lewi81,Apos10}. In Ref.\
\cite{Szmy13}, we gave an analytical proof of the result displayed in
Eq.\ (\ref{1.3}), and also we derived a closed-form formula for the
third-order derivative
$[\partial^{3}P_{\nu}(z)/\partial\nu^{3}]_{\nu=0}$. That work was
then extended by Laurenzi \cite{Laur15}, who found an expression for
the fourth-order derivative
$[\partial^{4}P_{\nu}(z)/\partial\nu^{4}]_{\nu=0}$.

The primary purpose of the present work is to pursue further the
research initiated by Schramkowski and continued by us
in Ref.\ \cite{Szmy13}. We shall show that for arbitrary
$n\in\mathbb{N}_{0}$ the second-order derivative
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$ may be expressed
in the form
\begin{equation}
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=n}
=-2P_{n}(z)\Li_{2}\frac{1-z}{2}+B_{n}(z)\ln\frac{z+1}{2}+C_{n}(z),
\label{1.5}
\end{equation}
where the polynomials $B_{n}(z)$ and $C_{n}(z)$ have the following
representations in terms of the Legendre polynomials:
\begin{equation}
B_{n}(z)=4[\psi(2n+1)-\psi(n+1)]P_{n}(z)
+4\sum_{k=0}^{n-1}\frac{2k+1}{(n-k)(n+k+1)}P_{k}(z)
\label{1.6}
\end{equation}
and
\begin{eqnarray}
C_{n}(z) &=& \left\{-\frac{\pi^{2}}{3}+4[\psi(2n+1)-\psi(n+1)]^{2}
+4\psi_{1}(2n+1)-2\psi_{1}(n+1)\right\}P_{n}(z)
\nonumber \\
&& +\,4\sum_{k=0}^{n-1}(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}
\bigg\{2\bigg[\psi(n+k+1)-\psi(n-k+1)
\nonumber \\
&& \qquad -\,\psi\left(\left\lfloor\frac{n+k}{2}\right\rfloor+1\right)
+\psi\left(\left\lfloor\frac{n-k}{2}\right\rfloor+1\right)\bigg]
\nonumber \\
&& \quad -\,(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}
-\frac{2n+1}{(n-k)(n+k+1)}\bigg\}P_{k}(z),
\label{1.7}
\end{eqnarray}
with $\psi(z)$ being already defined under Eq.\ (\ref{1.2}),
$\psi_{1}(z)=\mathrm{d}\psi(z)/\mathrm{d}z$, and $\lfloor
x\rfloor=\max\{n\in\mathbb{Z}:\:n\leqslant x\}$ standing for the
integer part of $x$. In fact, the above result for
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$, valid for
$n\in\mathbb{N}_{0}$, may be easily extended to any $n\in\mathbb{Z}$,
since with the use of the well-known identity
\begin{equation}
P_{-\nu-1}(z)=P_{\nu}(z)
\label{1.8}
\end{equation}
one immediately finds that
\begin{equation}
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=-n-1}
=\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=n}
\qquad (n\in\mathbb{N}_{0}).
\label{1.9}
\end{equation}

In addition to the above summarized study on the second-order
degree-derivative of the Legendre function of the first kind, which
will be presented in detail in Sec.\ \ref{II} below, later in Sec.\
\ref{III} we shall also prove that if $z\in\mathbb{C}\setminus[-1,1]$
and $n\in\mathbb{N}_{0}$, then the first-order derivative $[\partial
Q_{\nu}(z)/\partial\nu]_{\nu=n}$, where $Q_{\nu}(z)$ is the Legendre
function of the second kind, is given by
\begin{eqnarray}
\frac{\partial Q_{\nu}(z)}{\partial\nu}\bigg|_{\nu=n}
&=& -P_{n}(z)\Li_{2}\frac{1-z}{2}
-\frac{1}{2}P_{n}(z)\ln\frac{z+1}{2}\ln\frac{z-1}{2}
+\frac{1}{4}B_{n}(z)\ln\frac{z+1}{2}
\nonumber \\
&& -\,\frac{(-1)^{n}}{4}B_{n}(-z)\ln\frac{z-1}{2}
-\frac{\pi^{2}}{6}P_{n}(z)
+\frac{1}{4}C_{n}(z)-\frac{(-1)^{n}}{4}C_{n}(-z).
\label{1.10}
\end{eqnarray}
A counterpart expression for $[\partial
Q_{\nu}(x)/\partial\nu]_{\nu=n}$, applicable when $x\in(-1,1)$, will
also be derived.
%
%
\section{The derivatives
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$}
\label{II}
\setcounter{equation}{0}
\subsection{The general form of 
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$}
\label{II.1}
Our point of departure is the well-known recurrence relation
\begin{equation}
(\nu+1)P_{\nu+1}(z)-(2\nu+1)zP_{\nu}(z)+\nu P_{\nu-1}(z)=0
\label{2.1}
\end{equation}
obeyed by the Legendre functions of the first kind. Differentiating
it twice with respect to $\nu$ and setting then $\nu=n$ yields
\begin{eqnarray}
&& (n+1)\frac{\partial^{2}P_{\nu}(z)}
{\partial\nu^{2}}\bigg|_{\nu=n+1}
-(2n+1)z\frac{\partial^{2}P_{\nu}(z)}
{\partial\nu^{2}}\bigg|_{\nu=n}
+n\frac{\partial^{2}P_{\nu}(z)}
{\partial\nu^{2}}\bigg|_{\nu=n-1}
\nonumber \\
&& \qquad =\,-2\left[\frac{\partial P_{\nu}(z)}
{\partial\nu}\bigg|_{\nu=n+1}
-2z\frac{\partial P_{\nu}(z)}
{\partial\nu}\bigg|_{\nu=n}
+\frac{\partial P_{\nu}(z)}
{\partial\nu}\bigg|_{\nu=n-1}\right].
\label{2.2}
\end{eqnarray}
If we replace the first-order derivatives on the right-hand side with
expressions following from Eq.\ (\ref{1.1}), this furnishes
\begin{eqnarray}
&& (n+1)\frac{\partial^{2}P_{\nu}(z)}
{\partial\nu^{2}}\bigg|_{\nu=n+1}
-(2n+1)z\frac{\partial^{2}P_{\nu}(z)}
{\partial\nu^{2}}\bigg|_{\nu=n}
+n\frac{\partial^{2}P_{\nu}(z)}
{\partial\nu^{2}}\bigg|_{\nu=n-1}
\nonumber \\
&& \qquad =\,-2[P_{n+1}(z)-2zP_{n}(z)+P_{n-1}(z)]\ln\frac{z+1}{2}
-2[R_{n+1}(z)-2zR_{n}(z)+R_{n-1}(z)].
\nonumber \\
&&
\label{2.3}
\end{eqnarray}
From the formal point of view, one may look at Eq.\ (\ref{2.3}) as a
second-order difference equation and then two additional conditions
are necessary to single out the sequence
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$ from its general
solution. Such conditions may be chosen in a variety of ways but for
our purposes it is most convenient to take the explicit expression
for $[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=0}$, given in Eq.\
(\ref{1.1}), as the first one. The second suitable condition follows
from Eq.\ (\ref{2.3}) after one lets $n=0$. Then, with the use of the
identities
\begin{equation}
P_{-1}(z)=1, 
\qquad P_{0}(z)=1,
\qquad P_{1}(z)=z
\label{2.4}
\end{equation}
and \cite[Sec.\ 5.2]{Szmy06}
\begin{equation}
R_{-1}(z)=-2\ln\frac{z+1}{2},
\qquad R_{0}(z)=0,
\qquad R_{1}(z)=z-1,
\label{2.5}
\end{equation}
Eq.\ (\ref{2.3}) reduces to the form
\begin{equation}
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=1}
-z\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=0}
=2(z+1)\ln\frac{z+1}{2}-2(z-1).
\label{2.6}
\end{equation}
On combining Eq.\ (\ref{2.6}) with Eq.\ (\ref{1.1}), one finds that
\begin{equation}
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=1}
=-2z\Li_{2}\frac{1-z}{2}+2(z+1)\ln\frac{z+1}{2}-2(z-1).
\label{2.7}
\end{equation}
If necessary, Eq.\ (\ref{2.3}) may be applied recursively, with Eqs.\
(\ref{1.3}) and (\ref{2.7}) used as initial conditions, to generate
the derivative in question for any particular $n\geqslant2$. However,
as we shall show below, it is also possible to obtain a closed-form
representation for
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$.

As the first step towards that goal, we observe that the structure of
Eq.\ (\ref{2.3}), together with explicit expressions for the
derivatives $[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=0}$ and
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=1}$ displayed in Eqs.\
(\ref{1.3}) and (\ref{2.7}), fix the form of
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$ to be
\begin{equation}
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=n}
=A_{n}(z)\Li_{2}\frac{1-z}{2}+B_{n}(z)\ln\frac{z+1}{2}+C_{n}(z),
\label{2.8}
\end{equation}
where $A_{n}(z)$, $B_{n}(z)$ and $C_{n}(z)$ are polynomials in $z$ of
degree $n$. Since the right-hand side of Eq.\ (\ref{2.3}) does not
contain the dilogarithm function, the polynomial $A_{n}(z)$ solves
the homogeneous recurrence
\begin{equation}
(n+1)A_{n+1}(z)-(2n+1)zA_{n}(z)+nA_{n}(z)=0
\label{2.9}
\end{equation}
subject to the initial conditions
\begin{equation}
A_{0}(z)=-2=-2P_{0}(z),
\qquad
A_{1}(z)=-2z=-2P_{1}(z),
\label{2.10}
\end{equation}
which follow from Eqs.\ (\ref{1.1}) and (\ref{2.8}). Hence, we deduce
the following expression for $A_{n}(z)$ in terms of the Legendre
polynomial $P_{n}(z)$:
\begin{equation}
A_{n}(z)=-2P_{n}(z).
\label{2.11}
\end{equation}
Consequently, Eq.\ (\ref{2.8}) becomes
\begin{equation}
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=n}
=-2P_{n}(z)\Li_{2}\frac{1-z}{2}+B_{n}(z)\ln\frac{z+1}{2}+C_{n}(z).
\label{2.12}
\end{equation}
The representations of the polynomials $B_{n}(z)$ and $C_{n}(z)$
remain to be established.
%
%
\subsection{Differential equations for the polynomials $B_{n}(z)$ and
$C_{n}(z)$}
\label{II.2}
It is known that the Legendre function $P_{\nu}(z)$ obeys the
differential identity
\begin{equation}
\left[\frac{\mathrm{d}}{\mathrm{d}z}(1-z^{2})
\frac{\mathrm{d}}{\mathrm{d}z}+\nu(\nu+1)\right]P_{\nu}(z)=0.
\label{2.13}
\end{equation}
If we differentiate it twice with respect to $\nu$ and then put
$\nu=n$, this gives
\begin{equation}
\left[\frac{\mathrm{d}}{\mathrm{d}z}(1-z^{2})
\frac{\mathrm{d}}{\mathrm{d}z}+n(n+1)\right]
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=n}
=-2(2n+1)\frac{\partial P_{\nu}(z)}{\partial\nu}\bigg|_{\nu=n}
-2P_{n}(z)
\label{2.14}
\end{equation}
and further, after Eq.\ (\ref{1.1}) is plugged into the first term on
the right-hand side,
\begin{equation}
\left[\frac{\mathrm{d}}{\mathrm{d}z}(1-z^{2})
\frac{\mathrm{d}}{\mathrm{d}z}+n(n+1)\right]
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=n}
=-2(2n+1)P_{n}(z)\ln\frac{z+1}{2}-2(2n+1)R_{n}(z)-2P_{n}(z).
\label{2.15}
\end{equation}
Next, we insert Eq.\ (\ref{2.12}) into the left-hand side of Eq.\
(\ref{2.15}) and equate those terms appearing on both sides which
involve the logarithmic factor. This yields the following
inhomogeneous differential equation for $B_{n}(z)$:
\begin{equation}
\left[\frac{\mathrm{d}}{\mathrm{d}z}(1-z^{2})
\frac{\mathrm{d}}{\mathrm{d}z}+n(n+1)\right]B_{n}(z)
=4\left[(z+1)\frac{\mathrm{d}P_{n}(z)}{\mathrm{d}z}-nP_{n}(z)\right].
\label{2.16}
\end{equation}
Similarly, after equating polynomial expressions on both sides, we
arrive at the inhomogeneous equation for $C_{n}(z)$:
\begin{equation}
\left[\frac{\mathrm{d}}{\mathrm{d}z}(1-z^{2})
\frac{\mathrm{d}}{\mathrm{d}z}+n(n+1)\right]C_{n}(z)
=2(z-1)\frac{\mathrm{d}B_{n}(z)}{\mathrm{d}z}+B_{n}(z)
-2(2n+1)R_{n}(z).
\label{2.17}
\end{equation}

Consider Eq.\ (\ref{2.16}). It is evident that it does not possess a
unique polynomial solution, since to any particular solution of that
form one may add an arbitrary multiple of the Legendre polynomial
$P_{n}(z)$, which results in another polynomial solution. To make the
polynomial solution unique, we thus need an additional constraint.
The latter follows from the limiting relation \cite{Magn66}
\begin{equation}
P_{\nu}(z)\stackrel{z\to-1}{\longrightarrow}
\frac{\sin(\pi\nu)}{\pi}\ln\frac{z+1}{2}+O(1),
\label{2.18}
\end{equation}
from which we find
\begin{equation}
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=n}
\stackrel{z\to-1}{\longrightarrow}O(1).
\label{2.19}
\end{equation}
Hence, the left-hand side of Eq.\ (\ref{2.12}) remains finite for
$z\to-1$ and to make the right-hand side also finite in that limit,
we are forced to put
\begin{equation}
B_{n}(-1)=0.
\label{2.20}
\end{equation}
If, in turn, we wish to make the polynomial solution to Eq.\
(\ref{2.17}) unique, we use the identity
\begin{equation}
P_{\nu}(1)=1.
\label{2.21}
\end{equation}
Differentiating twice with respect to $\nu$, we obtain
\begin{equation}
\frac{\partial^{2}P_{\nu}(1)}{\partial\nu^{2}}\bigg|_{\nu=n}=0.
\label{2.22} 
\end{equation}
Since $\Li_{2}0=0$ and $\ln1=0$, we deduce that $C_{n}(z)$ is
constrained to obey
\begin{equation}
C_{n}(1)=0.
\label{2.23}
\end{equation}
Below we shall exploit Eqs.\ (\ref{2.16}), (\ref{2.20}), (\ref{2.17})
and (\ref{2.23}) to determine the polynomials $B_{n}(z)$ and
$C_{n}(z)$.
%
%
\subsection{Construction of the polynomials $B_{n}(z)$}
\label{II.3}
A general form of the polynomials $B_{n}(z)$ may be obtained with
ease. In Ref.\ \cite[Sec.\ 5.2.2]{Szmy06}, we have found that the
polynomials $R_{n}(z)$ obey the differential relation
\begin{equation}
\left[\frac{\mathrm{d}}{\mathrm{d}z}(1-z^{2})
\frac{\mathrm{d}}{\mathrm{d}z}+n(n+1)\right]R_{n}(z)
=2\left[(z-1)\frac{\mathrm{d}P_{n}(z)}{\mathrm{d}z}-nP_{n}(z)\right].
\label{2.24}
\end{equation}
Hence, with the use of the well-known identity
\begin{equation}
P_{n}(-z)=(-1)^{n}P_{n}(z),
\label{2.25}
\end{equation}
we deduce that
\begin{equation}
\left[\frac{\mathrm{d}}{\mathrm{d}z}(1-z^{2})
\frac{\mathrm{d}}{\mathrm{d}z}+n(n+1)\right]R_{n}(-z)
=(-1)^{n}2\left[(z+1)\frac{\mathrm{d}P_{n}(z)}{\mathrm{d}z}
-nP_{n}(z)\right].
\label{2.26}
\end{equation}
Comparison of Eqs.\ (\ref{2.16}) and (\ref{2.26}) shows that the
polynomial $B_{n}(z)$ must be of the form
\begin{equation}
B_{n}(z)=(-1)^{n}2R_{n}(-z)+b_{n}P_{n}(z).
\label{2.27}
\end{equation}
To determine the constant $b_{n}$, we put $z=-1$ in Eq.\
(\ref{2.27}). By virtue of the constraint (\ref{2.20}), with the use
of the relations
\begin{equation}
P_{n}(-1)=(-1)^{n}
\label{2.28}
\end{equation}
and (cf.\ Ref.\ \cite[Eq.\ (5.10)]{Szmy06})
\begin{equation}
R_{n}(1)=0,
\label{2.29}
\end{equation}
we infer that
\begin{equation}
b_{n}=0,
\label{2.30}
\end{equation}
and thus finally we arrive at
\begin{equation}
B_{n}(z)=(-1)^{n}2R_{n}(-z).
\label{2.31}
\end{equation}
On combining Eq.\ (\ref{2.31}) with Eq.\ (\ref{1.2}), we have the
following explicit representation of $B_{n}(z)$:
\begin{equation}
B_{n}(z)=4[\psi(2n+1)-\psi(n+1)]P_{n}(z)
+4\sum_{k=0}^{n-1}\frac{2k+1}{(n-k)(n+k+1)}P_{k}(z).
\label{2.32}
\end{equation}
Further expressions for $B_{n}(z)$ may be obtained if one combines
Eq.\ (\ref{2.31}) with Eqs.\ (2.1), (2.2), (2.4) and (5.90) from
Ref.\ \cite{Szmy06} or with Eqs.\ (11) and (12) from Ref.\
\cite{Szmy07}.
%
%
\subsection{Construction of the polynomials $C_{n}(z)$}
\label{II.4}
We shall seek a representation of $C_{n}(z)$ in the form of a linear
combination of Legendre polynomials:
\begin{equation}
C_{n}(z)=\sum_{k=0}^{n}c_{nk}P_{k}(z).
\label{2.33}
\end{equation}
Action on both sides of Eq.\ (\ref{2.33}) with the Legendre
differential operator appearing on the left-hand side of Eq.\
(\ref{2.17}) gives
\begin{equation}
\left[\frac{\mathrm{d}}{\mathrm{d}z}(1-z^{2})
\frac{\mathrm{d}}{\mathrm{d}z}+n(n+1)\right]C_{n}(z)
=\sum_{k=0}^{n-1}(n-k)(n+k+1)c_{nk}P_{k}(z).
\label{2.34}
\end{equation}
On the other side, with the aid of Eqs.\ (\ref{1.2}) and
(\ref{2.32}), and of the identity
\begin{equation}
(z-1)\frac{\mathrm{d}P_{n}(z)}{\mathrm{d}z}
=nP_{n}(z)+\sum_{k=0}^{n-1}(-1)^{n+k}(2k+1)P_{k}(z),
\label{2.35}
\end{equation}
after some algebra we find that the expression on the right-hand side
of Eq.\ (\ref{2.17}) may be written as
\begin{eqnarray}
&& \hspace*{-3em}
2(z-1)\frac{\mathrm{d}B_{n}(z)}{\mathrm{d}z}+B_{n}(z)-2(2n+1)R_{n}(z)
\nonumber \\
&=& \sum_{k=0}^{n-1}\bigg\{(-1)^{n+k}8(2k+1)[\psi(2n+1)-\psi(n+1)]
\nonumber \\
&& +\,8(2k+1)\sum_{m=k}^{n-1}(-1)^{m+k}\frac{2m+1}{(n-m)(n+m+1)}
\nonumber \\
&& -\,\frac{4(2k+1)^{2}}{(n-k)(n+k+1)}
-(-1)^{n+k}\frac{4(2n+1)(2k+1)}{(n-k)(n+k+1)}\bigg\}P_{k}(z).
\label{2.36}
\end{eqnarray}
Equations (\ref{2.17}), (\ref{2.34}) and (\ref{2.36}) yield
\begin{eqnarray}
c_{nk} &=& (-1)^{n+k}[\psi(2n+1)-\psi(n+1)]
\frac{8(2k+1)}{(n-k)(n+k+1)}
\nonumber \\
&& +\,\frac{8(2k+1)}{(n-k)(n+k+1)}
\sum_{m=k}^{n-1}(-1)^{m+k}\frac{2m+1}{(n-m)(n+m+1)}
\nonumber \\
&& -\,\frac{4(2k+1)^{2}}{(n-k)^{2}(n+k+1)^{2}}
-(-1)^{n+k}\frac{4(2n+1)(2k+1)}{(n-k)^{2}(n+k+1)^{2}}
\qquad (0\leqslant k\leqslant n-1).
\nonumber \\
&&
\label{2.37}
\end{eqnarray}
It is proven in Appendix \ref{A.I} that
\begin{eqnarray}
\sum_{m=k}^{n-1}(-1)^{n+m}\frac{2m+1}{(n-m)(n+m+1)}
&=& -\psi(2n+1)+\psi(n+k+1)+\psi(n+1)-\psi(n-k+1)
\nonumber \\
&& -\,\psi\left(\left\lfloor\frac{n+k}{2}\right\rfloor+1\right)
+\psi\left(\left\lfloor\frac{n-k}{2}\right\rfloor+1\right)
\nonumber \\
&& \hspace*{10em} (0\leqslant k\leqslant n-1),
\label{2.38}
\end{eqnarray}
where $\lfloor x\rfloor=\max\{n\in\mathbb{Z}:\:n\leqslant x\}$. Use
of Eq.\ (\ref{2.38}) casts Eq.\ (\ref{2.37}) into the final form
\begin{eqnarray}
c_{nk} &=& (-1)^{n+k}\frac{8(2k+1)}{(n-k)(n+k+1)}
\nonumber \\
&& \quad \times\bigg[\psi(n+k+1)-\psi(n-k+1)
-\psi\left(\left\lfloor\frac{n+k}{2}\right\rfloor+1\right)
+\psi\left(\left\lfloor\frac{n-k}{2}\right\rfloor+1\right)\bigg]
\nonumber \\
&& -\,\frac{4(2k+1)^{2}}{(n-k)^{2}(n+k+1)^{2}}
-(-1)^{n+k}\frac{4(2n+1)(2k+1)}{(n-k)^{2}(n+k+1)^{2}}
\qquad (0\leqslant k\leqslant n-1).
\nonumber \\
&&
\label{2.39}
\end{eqnarray}

Equation (\ref{2.39}) says nothing about the coefficient $c_{nn}$.
But from Eqs.\ (\ref{2.23}), (\ref{2.21}) and (\ref{2.33}) it can be
deduced that $c_{nn}$ may be expressed as
\begin{equation}
c_{nn}=-\sum_{k=0}^{n-1}c_{nk}.
\label{2.40}
\end{equation}
This implies that the polynomial $C_{n}(z)$ may be written as
\begin{equation}
C_{n}(z)=\sum_{k=0}^{n-1}c_{nk}[P_{k}(z)-P_{n}(z)],
\label{2.41}
\end{equation}
or explicitly, if the result in Eq.\ (\ref{2.39}) is used, as
\begin{eqnarray}
C_{n}(z) &=& 4\sum_{k=0}^{n-1}(-1)^{n+k}
\frac{2k+1}{(n-k)(n+k+1)}
\nonumber \\
&& \times\,\bigg\{2\bigg[\psi(n+k+1)-\psi(n-k+1)
-\psi\left(\left\lfloor\frac{n+k}{2}\right\rfloor+1\right)
+\psi\left(\left\lfloor\frac{n-k}{2}\right\rfloor+1\right)\bigg]
\nonumber \\
&& \quad -\,(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}
-\frac{2n+1}{(n-k)(n+k+1)}\bigg\}[P_{k}(z)-P_{n}(z)].
\label{2.42}
\end{eqnarray}

A bit different formula for $C_{n}(z)$ is obtained if the coefficient
$c_{nn}$ is expressed in a closed form. To find the latter, we
combine Eqs.\ (\ref{2.37}) and (\ref{2.40}) and write
\begin{eqnarray}
c_{nn} &=& -8[\psi(2n+1)-\psi(n+1)]
\sum_{k=0}^{n-1}(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}
\nonumber \\
&& -\,8\sum_{k=0}^{n-1}(-1)^{k}\frac{2k+1}{(n-k)(n+k+1)}
\sum_{m=k}^{n-1}(-1)^{m}\frac{2m+1}{(n-m)(n+m+1)}
\nonumber \\
&& +\,4\sum_{k=0}^{n-1}\frac{(2k+1)^{2}}{(n-k)^{2}(n+k+1)^{2}}
+4\sum_{k=0}^{n-1}(-1)^{n+k}
\frac{(2n+1)(2k+1)}{(n-k)^{2}(n+k+1)^{2}}.
\label{2.43}
\end{eqnarray}
The sums appearing in Eq.\ (\ref{2.43}) are evaluated in individual
subsections of the appendix, where it is found that
\begin{eqnarray}
&& \sum_{k=0}^{n-1}(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}
=-\psi(2n+1)+\psi(n+1),
\label{2.44}
\end{eqnarray}
\begin{eqnarray}
&& \hspace*{-5em}
\sum_{k=0}^{n-1}(-1)^{k}\frac{2k+1}{(n-k)(n+k+1)}
\sum_{m=k}^{n-1}(-1)^{m}\frac{2m+1}{(n-m)(n+m+1)}
\nonumber \\
&& =\,\frac{\pi^{2}}{12}-\frac{\gamma}{2n+1}
+\frac{1}{2}[\psi(2n+1)-\psi(n+1)]^{2}
-\frac{1}{2n+1}\psi(2n+1)-\frac{1}{2}\psi_{1}(2n+1),
\nonumber \\
&&
\label{2.45}
\end{eqnarray}
\begin{equation}
\sum_{k=0}^{n-1}\frac{(2k+1)^{2}}{(n-k)^{2}(n+k+1)^{2}}
=\frac{\pi^{2}}{6}-\frac{2\gamma}{2n+1}-\frac{2}{2n+1}\psi(2n+1)
-\psi_{1}(2n+1)
\label{2.46}
\end{equation}
and
\begin{equation}
\sum_{k=0}^{n-1}(-1)^{n+k}\frac{(2n+1)(2k+1)}{(n-k)^{2}(n+k+1)^{2}}
=-\frac{\pi^{2}}{12}+\psi_{1}(2n+1)-\frac{1}{2}\psi_{1}(n+1),
\label{2.47}
\end{equation}
with $\gamma$ standing for the Euler--Mascheroni constant and with
$\psi_{1}(\zeta)=\mathrm{d}\psi(\zeta)/\mathrm{d}\zeta$ being the
trigamma function. Plugging the results (\ref{2.44})--(\ref{2.47})
into the right-hand side of Eq.\ (\ref{2.43}) furnishes the
coefficient $c_{nn}$ in the compact form
\begin{equation}
c_{nn}=-\frac{\pi^{2}}{3}+4[\psi(2n+1)-\psi(n+1)]^{2}
+4\psi_{1}(2n+1)-2\psi_{1}(n+1).
\label{2.48}
\end{equation}
Hence, by virtue of Eqs.\ (\ref{2.33}), (\ref{2.39}) and
(\ref{2.48}), we eventually arrive at the sought formula
\begin{eqnarray}
C_{n}(z) &=& \left\{-\frac{\pi^{2}}{3}+4[\psi(2n+1)-\psi(n+1)]^{2}
+4\psi_{1}(2n+1)-2\psi_{1}(n+1)\right\}P_{n}(z)
\nonumber \\
&& +\,4\sum_{k=0}^{n-1}(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}
\bigg\{2\bigg[\psi(n+k+1)-\psi(n-k+1)
\nonumber \\
&& \qquad -\,\psi\left(\left\lfloor\frac{n+k}{2}\right\rfloor+1\right)
+\psi\left(\left\lfloor\frac{n-k}{2}\right\rfloor+1\right)\bigg]
\nonumber \\
&& \quad -\,(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}
-\frac{2n+1}{(n-k)(n+k+1)}\bigg\}P_{k}(z),
\label{2.49}
\end{eqnarray}
alternative to the one in Eq.\ (\ref{2.42}).
%
%
\subsection{Explicit expressions for 
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$ with $0\leqslant
n\leqslant3$}
\label{II.5}
It may be of interest to see how the derivatives
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$ look explicitly
for several lowest values of $n$. From Eqs.\ (\ref{2.12}),
(\ref{2.32}) and (\ref{2.49}), for $0\leqslant n\leqslant3$ we find
that
\begin{subequations}
\begin{eqnarray}
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=0}
&=& -2\Li_{2}\frac{1-z}{2}, 
\label{2.50a}
\\*[1ex]
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=1}
&=& -2z\Li_{2}\frac{1-z}{2}+(2z+2)\ln\frac{z+1}{2}-2z+2,
\label{2.50b}
\\*[1ex]
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=2}
&=& (-3z^{2}+1)\Li_{2}\frac{1-z}{2}
+\left(\frac{7}{2}z^{2}+3z-\frac{1}{2}\right)\ln\frac{z+1}{2}
-\frac{11}{4}z^{2}+\frac{5}{2}z+\frac{1}{4},
\nonumber \\
&&
\label{2.50c}
\\*[1ex]
\frac{\partial^{2}P_{\nu}(z)}{\partial\nu^{2}}\bigg|_{\nu=3}
&=& (-5z^{3}+3z)\Li_{2}\frac{1-z}{2}
+\left(\frac{37}{6}z^{3}+5z^{2}-\frac{5}{2}z
-\frac{4}{3}\right)\ln\frac{z+1}{2}
\nonumber \\
&& -\,\frac{155}{36}z^{3}+\frac{23}{6}z^{2}
+\frac{19}{12}z-\frac{10}{9}.
\label{2.50d}
\end{eqnarray}
\label{2.50}%
\end{subequations}
%
%
\section{The derivatives $[\partial Q_{\nu}(z)/\partial\nu]_{\nu=n}$
and $[\partial Q_{\nu}(x)/\partial\nu]_{\nu=n}$}
\label{III}
\setcounter{equation}{0}
In Refs.\ \cite{Szmy06,Szmy07}, we exploited representations of the
first-order derivatives $[\partial P_{\nu}(z)/\partial\nu]_{\nu=n}$
found therein to obtain expressions for the Legendre functions of the
second kind $Q_{n}(z)$, with $n\in\mathbb{N}_{0}$, both for
$z\in\mathbb{C}\setminus[-1,1]$ and for $z=x\in(-1,1)$. Below we
shall show that the knowledge of the second-order derivatives
$[\partial^{2}P_{\nu}(z)/\partial\nu^{2}]_{\nu=n}$ allows one to
obtain explicit formulas for the first-order derivatives $[\partial
Q_{\nu}(z)/\partial\nu]_{\nu=n}$ and $[\partial
Q_{\nu}(x)/\partial\nu]_{\nu=n}$, again with $n\in\mathbb{N}_{0}$.
\subsection{The derivatives $[\partial
Q_{\nu}(z)/\partial\nu]_{\nu=n}$ for $z\in\mathbb{C}\setminus[-1,1]$}
\label{III.1}
\subsubsection{The general form of $[\partial
Q_{\nu}(z)/\partial\nu]_{\nu=n}$}
\label{III.1.1}
The Legendre function of the second kind, $Q_{\nu}(z)$, may be
defined in terms of the function of the first kind through the
formula
\begin{equation}
Q_{\nu}(z)=\frac{\pi}{2}
\frac{\mathrm{e}^{\mp\mathrm{i}\pi\nu}P_{\nu}(z)-P_{\nu}(-z)}
{\sin(\pi\nu)}
\qquad (\Imag z\gtrless0).
\label{3.1}
\end{equation}
Hence, it follows that
\begin{eqnarray}
&& \frac{\partial
Q_{\nu}(z)}{\partial\nu}=\frac{\pi}{2\sin^{2}(\pi\nu)}
\left\{-\pi[P_{\nu}(z)-P_{\nu}(-z)\cos(\pi\nu)]
+\left[\mathrm{e}^{\mp\mathrm{i}\pi\nu}
\frac{\partial P_{\nu}(z)}{\partial\nu}
-\frac{\partial P_{\nu}(-z)}{\partial\nu}\right]\sin(\pi\nu)\right\}
\nonumber \\
&& \hspace*{30em} (\Imag z\gtrless0).
\label{3.2}
\end{eqnarray}
From this, for $\nu=n\in\mathbb{N}_{0}$, with the use of the
L'Hospital rule, we obtain
\begin{equation}
\frac{\partial Q_{\nu}(z)}{\partial\nu}\bigg|_{\nu=n}
=-\frac{\pi^{2}}{4}P_{n}(z)\mp\frac{\mathrm{i}\pi}{2}
\frac{\partial P_{\nu}(z)}{\partial\nu}\bigg|_{\nu=n}
+\frac{1}{4}\frac{\partial^{2}P_{\nu}(z)}
{\partial\nu^{2}}\bigg|_{\nu=n}
-\frac{(-1)^{n}}{4}\frac{\partial^{2}P_{\nu}(-z)}{\partial\nu^{2}}
\bigg|_{\nu=n}
\qquad (\Imag z\gtrless0).
\label{3.3}
\end{equation}
If in the above formula the second-order derivatives
$[\partial^{2}P_{\nu}(\pm z)/\partial\nu^{2}]_{\nu=n}$ are
substituted with expressions following from Eq.\ (\ref{2.12}) and the
first-order derivative $[\partial P_{\nu}(z)/\partial\nu]_{\nu=n}$ is
replaced by the right-hand side of Eq.\ (\ref{1.1}), this yields
$[\partial Q_{\nu}(z)/\partial\nu]_{\nu=n}$ in the form
\begin{eqnarray}
\frac{\partial Q_{\nu}(z)}{\partial\nu}\bigg|_{\nu=n}
&=& \frac{1}{2}P_{n}(z)
\left(\Li_{2}\frac{z+1}{2}-\Li_{2}\frac{1-z}{2}\right)
+\left[\frac{1}{4}B_{n}(z)\mp\frac{\mathrm{i}\pi}{2}P_{n}(z)\right]
\ln\frac{z+1}{2}
\nonumber \\
&& -\,\frac{(-1)^{n}}{4}B_{n}(-z)\ln\frac{1-z}{2}
-\frac{\pi^{2}}{4}P_{n}(z)
\mp\frac{\mathrm{i}\pi}{2}R_{n}(z)+\frac{1}{4}C_{n}(z)
-\frac{(-1)^{n}}{4}C_{n}(-z)
\nonumber \\
&& \hspace*{25em} (\Imag z\gtrless0).
\label{3.4}
\end{eqnarray}
A more elegant expression for $[\partial
Q_{\nu}(z)/\partial\nu]_{\nu=n}$ follows if the dilogarithm
$\Li_{2}[(z+1)/2]$ is eliminated from Eq.\ (\ref{3.4}) with the aid
of the Euler's identity \cite[Eq.\ (1.11)]{Lewi81}
\begin{equation}
\Li_{2}z+\Li_{2}(1-z)=\frac{\pi^{2}}{6}-\ln z\ln(1-z),
\label{3.5}
\end{equation}
the relation
\begin{equation}
1-z=\mathrm{e}^{\mp\mathrm{i}\pi}(z-1)
\qquad (\Imag z\gtrless0)
\label{3.6}
\end{equation}
and the result in Eq.\ (\ref{2.31}). Proceeding in that way, one
eventually finds that
\begin{eqnarray}
\frac{\partial Q_{\nu}(z)}{\partial\nu}\bigg|_{\nu=n}
&=& -P_{n}(z)\Li_{2}\frac{1-z}{2}
-\frac{1}{2}P_{n}(z)\ln\frac{z+1}{2}\ln\frac{z-1}{2}
+\frac{1}{4}B_{n}(z)\ln\frac{z+1}{2}
\nonumber \\
&& -\,\frac{(-1)^{n}}{4}B_{n}(-z)\ln\frac{z-1}{2}
-\frac{\pi^{2}}{6}P_{n}(z)
+\frac{1}{4}C_{n}(z)-\frac{(-1)^{n}}{4}C_{n}(-z)
\qquad (n\in\mathbb{N}_{0}).
\nonumber \\
&&
\label{3.7}
\end{eqnarray}
\subsubsection{Explicit expressions for $[\partial
Q_{\nu}(z)/\partial\nu]_{\nu=n}$ with $0\leqslant n\leqslant3$}
\label{III.1.2}
Explicit forms of the derivatives $[\partial
Q_{\nu}(z)/\partial\nu]_{\nu=n}$ with $0\leqslant n\leqslant3$,
obtained from Eq.\ (\ref{3.7}) with the use of Eqs.\ (\ref{2.32}) and
(\ref{2.49}), are
\begin{subequations}
\begin{eqnarray}
\frac{\partial Q_{\nu}(z)}{\partial\nu}\bigg|_{\nu=0}
&=& -\Li_{2}\frac{1-z}{2}-\frac{1}{2}\ln\frac{z+1}{2}\ln\frac{z-1}{2}
-\frac{\pi^{2}}{6},
\label{3.8a}
\\*[1ex]
\frac{\partial Q_{\nu}(z)}{\partial\nu}\bigg|_{\nu=1}
&=& -z\Li_{2}\frac{1-z}{2}
-\frac{1}{2}z\ln\frac{z+1}{2}\ln\frac{z-1}{2}
\nonumber \\
&& +\,\left(\frac{1}{2}z+\frac{1}{2}\right)\ln\frac{z+1}{2}
+\left(-\frac{1}{2}z+\frac{1}{2}\right)\ln\frac{z-1}{2}
-\frac{\pi^{2}}{6}z+1,
\label{3.8b}
\\*[1ex]
\frac{\partial Q_{\nu}(z)}{\partial\nu}\bigg|_{\nu=2}
&=& \left(-\frac{3}{2}z^{2}+\frac{1}{2}\right)\Li_{2}\frac{1-z}{2}
+\left(-\frac{3}{4}z^{2}+\frac{1}{4}\right)
\ln\frac{z+1}{2}\ln\frac{z-1}{2}
\nonumber \\
&& +\,\left(\frac{7}{8}z^{2}+\frac{3}{4}z-\frac{1}{8}\right)
\ln\frac{z+1}{2}
+\left(-\frac{7}{8}z^{2}+\frac{3}{4}z+\frac{1}{8}\right)
\ln\frac{z-1}{2}-\frac{\pi^{2}}{4}z^{2}
+\frac{5}{4}z+\frac{\pi^{2}}{12},
\nonumber \\
&&
\label{3.8c}
\\*[1ex]
\frac{\partial Q_{\nu}(z)}{\partial\nu}\bigg|_{\nu=3}
&=& \left(-\frac{5}{2}z^{3}+\frac{3}{2}z\right)\Li_{2}\frac{1-z}{2}
+\left(-\frac{5}{4}z^{3}+\frac{3}{4}z\right)
\ln\frac{z+1}{2}\ln\frac{z-1}{2}
\nonumber \\
&& +\,\left(\frac{37}{24}z^{3}+\frac{5}{4}z^{2}
-\frac{5}{8}z-\frac{1}{3}\right)\ln\frac{z+1}{2}
+\left(-\frac{37}{24}z^{3}+\frac{5}{4}z^{2}
+\frac{5}{8}z-\frac{1}{3}\right)\ln\frac{z-1}{2}
\nonumber \\
&& -\,\frac{5\pi^{2}}{12}z^{3}+\frac{23}{12}z^{2}
+\frac{\pi^{2}}{4}z-\frac{5}{9}.
\label{3.8d}
\end{eqnarray}
\label{3.8}%
\end{subequations}
We find it remarkable that coefficients in the polynomial part of
$[\partial Q_{\nu}(z)/\partial\nu]_{\nu=n}$ are alternately
irrational and rational.
\subsection{The derivatives $[\partial
Q_{\nu}(x)/\partial\nu]_{\nu=n}$ for $-1<x<1$}
\label{III.2}
On the real interval $-1<x<1$, the Legendre function of the second
kind, $Q_{\nu}(x)$, is defined as the average of the limits
$Q_{\nu}(x+\mathrm{i}0)$ and $Q_{\nu}(x-\mathrm{i}0)$ resulting when
$z$ approaches $x$ from the upper ($\Imag z>0$) and lower ($\Imag
z<0$) half-planes, respectively. One has
\begin{equation}
Q_{\nu}(x)=\frac{1}{2}[Q_{\nu}(x+\mathrm{i}0)+Q_{\nu}(x-\mathrm{i}0)]
\qquad (-1<x<1),
\label{3.9}
\end{equation}
and consequently
\begin{equation}
\frac{\partial Q_{\nu}(x)}{\partial\nu}\bigg|_{\nu=n}
=\frac{1}{2}\frac{\partial Q_{\nu}(x+\mathrm{i}0)}
{\partial\nu}\bigg|_{\nu=n}
+\frac{1}{2}\frac{\partial Q_{\nu}(x-\mathrm{i}0)}
{\partial\nu}\bigg|_{\nu=n}.
\label{3.10}
\end{equation}
From this, with the use of Eq.\ (\ref{3.7}) and the identity
\begin{equation}
x-1\pm\mathrm{i}0=\mathrm{e}^{\pm\mathrm{i}\pi}(1-x)
\qquad (-1<x<1),
\label{3.11}
\end{equation}
one finds that
\begin{eqnarray}
\frac{\partial Q_{\nu}(x)}{\partial\nu}\bigg|_{\nu=n}
&=& -P_{n}(x)\Li_{2}\frac{1-x}{2}
-\frac{1}{2}P_{n}(x)\ln\frac{1+x}{2}\ln\frac{1-x}{2}
+\frac{1}{4}B_{n}(x)\ln\frac{1+x}{2}
\nonumber \\
&& -\,\frac{(-1)^{n}}{4}B_{n}(-x)\ln\frac{1-x}{2}
-\frac{\pi^{2}}{6}P_{n}(x)
+\frac{1}{4}C_{n}(x)-\frac{(-1)^{n}}{4}C_{n}(-x)
\qquad (n\in\mathbb{N}_{0}).
\nonumber \\
&&
\label{3.12}
\end{eqnarray}

There is no need to provide here explicit representations for the
derivatives $[\partial Q_{\nu}(x)/\partial\nu]_{\nu=n}$ for several
lowest non-negative values of $n$. As it is seen from Eqs.\
(\ref{3.7}) and (\ref{3.12}), such representations for $0\leqslant
n\leqslant3$ may be immediately deduced from Eqs.\
(\ref{3.8a})--(\ref{3.8d}) after the replacement of $z$ with $x$ is
made everywhere in the latter set of equations, except for the
logarithm $\ln[(z-1)/2]$, which is to be substituted with
$\ln[(1-x)/2]$.
\section*{Acknowledgments}
I wish to thank Dr.\ George P.\ Schramkowski for kindly communicating
to me the formula in Eq.\ (\ref{1.3}) and for the subsequent
inspiring correspondence.
%
%
\appendix
\renewcommand{\theequation}{\thesubsection.\arabic{equation}}
\section{Appendix: Proofs of summation formulas used in Sec.\
\ref{II.4}} 
\label{A} 
\setcounter{equation}{0}
\subsection{The summation formulas (\ref{2.38}) and (\ref{2.44})}
\label{A.I} 
\setcounter{equation}{0}
We denote
\begin{equation}
S_{1}=\sum_{k=m}^{n-1}(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}
\qquad (0\leqslant m\leqslant n-1).
\label{A.1}
\end{equation}
We have
\begin{equation}
S_{1}=\sum_{k=m}^{n-1}\frac{(-1)^{n+k}}{n-k}
-\sum_{k=m}^{n-1}\frac{(-1)^{n+k}}{n+k+1}
\label{A.2}
\end{equation}
and further
\begin{equation}
S_{1}=\sum_{k=1}^{n-m}\frac{(-1)^{k}}{k}
+\sum_{k=n+m+1}^{2n}\frac{(-1)^{k}}{k}
=\sum_{k=1}^{n-m}\frac{(-1)^{k}}{k}
+\sum_{k=1}^{2n}\frac{(-1)^{k}}{k}
-\sum_{k=1}^{n+m}\frac{(-1)^{k}}{k}.
\label{A.3}
\end{equation}
However, it is easy to show that
\begin{equation}
\sum_{k=1}^{N}\frac{(-1)^{k}}{k}
=-\sum_{k=1}^{N}\frac{1}{k}
+\sum_{k=1}^{\lfloor N/2\rfloor}\frac{1}{k}
\qquad (N\in\mathbb{N}_{0}),
\label{A.4}
\end{equation}
where $\lfloor x\rfloor=\max\{n\in\mathbb{Z}:\:n\leqslant x\}$ stands
for the integer part of $x$. Since
\begin{equation}
\sum_{k=1}^{N}\frac{1}{k}
=\psi(N+1)-\psi(1)
\qquad (N\in\mathbb{N}_{0}),
\label{A.5}
\end{equation}
with $\psi(z)=\mathrm{d}\ln\Gamma(z)/\mathrm{d}z$ being the digamma
function, Eq.\ (\ref{A.4}) may be rewritten in the form
\begin{equation}
\sum_{k=1}^{N}\frac{(-1)^{k}}{k}=-\psi(N+1)
+\psi\left(\bigg\lfloor\frac{N}{2}\bigg\rfloor+1\right)
\qquad (N\in\mathbb{N}_{0}).
\label{A.6}
\end{equation}
Application of the result (\ref{A.6}) to each of the three sums on
the extreme right-hand side of Eq.\ (\ref{A.3}) gives finally
\begin{eqnarray}
\sum_{k=m}^{n-1}(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}
&=& -\psi(2n+1)+\psi(n+m+1)+\psi(n+1)-\psi(n-m+1)
\nonumber \\
&& -\,\psi\left(\bigg\lfloor\frac{n+m}{2}\bigg\rfloor+1\right)
+\psi\left(\bigg\lfloor\frac{n-m}{2}\bigg\rfloor+1\right)
\nonumber \\
&& \hspace*{13em} (0\leqslant m\leqslant n-1).
\label{A.7}
\end{eqnarray}
After $k$ is interchanged with $m$, Eq.\ (\ref{A.7}) becomes
identical with Eq.\ (\ref{2.38}).

For $m=0$, Eq.\ (\ref{A.7}) becomes
\begin{equation}
\sum_{k=0}^{n-1}(-1)^{n+k}\frac{2k+1}{(n-k)(n+k+1)}
=-\psi(2n+1)+\psi(n+1),
\label{A.8}
\end{equation}
which is Eq.\ (\ref{2.44}).
%
%
\subsection{The summation formula (\ref{2.45})}
\label{A.II}
\setcounter{equation}{0}
We denote
\begin{equation}
S_{2}=\sum_{k=0}^{n-1}(-1)^{k}\frac{2k+1}{(n-k)(n+k+1)}
\sum_{m=k}^{n-1}(-1)^{m}\frac{2m+1}{(n-m)(n+m+1)}.
\label{A.9}
\end{equation}
Application of the identity
\begin{equation}
\sum_{k=N_{1}}^{N_{2}}f_{k}\sum_{m=k}^{N_{2}}f_{m}
=\frac{1}{2}\left(\sum_{k=N_{1}}^{N_{2}}f_{k}\right)^{2}
+\frac{1}{2}\sum_{k=N_{1}}^{N_{2}}f_{k}^{2}
\qquad (N_{1}\leqslant N_{2})
\label{A.10}
\end{equation}
transforms Eq.\ (\ref{A.9}) into
\begin{equation}
S_{2}=\frac{1}{2}\left[\sum_{k=0}^{n-1}(-1)^{k}
\frac{2k+1}{(n-k)(n+k+1)}\right]^{2}
+\frac{1}{2}\sum_{k=0}^{n-1}\frac{(2k+1)^{2}}{(n-k)^{2}(n+k+1)^{2}},
\label{A.11}
\end{equation}
from which, with the help of Eqs.\ (\ref{A.8}) and (\ref{A.22}),
we obtain
\begin{eqnarray}
&& \hspace*{-5em}
\sum_{k=0}^{n-1}(-1)^{k}\frac{2k+1}{(n-k)(n+k+1)}
\sum_{m=k}^{n-1}(-1)^{m}\frac{2m+1}{(n-m)(n+m+1)}
\nonumber \\
&& =\,\frac{\pi^{2}}{12}-\frac{\gamma}{2n+1}
+\frac{1}{2}[\psi(2n+1)-\psi(n+1)]^{2}
-\frac{1}{2n+1}\psi(2n+1)-\frac{1}{2}\psi_{1}(2n+1),
\nonumber \\
&&
\label{A.12}
\end{eqnarray}
which is Eq.\ (\ref{2.45}).

To prove the identity (\ref{A.10}), we write the obvious chain of
equalities ($N_{1}\leqslant N_{2}$ is assumed)
\begin{equation}
\left(\sum_{k=N_{1}}^{N_{2}}f_{k}\right)^{2}
=\sum_{k=N_{1}}^{N_{2}}f_{k}\sum_{m=N_{1}}^{N_{2}}f_{m}
=\sum_{k=N_{1}}^{N_{2}}f_{k}\sum_{m=N_{1}}^{k}f_{m}
+\sum_{k=N_{1}}^{N_{2}}f_{k}\sum_{m=k}^{N_{2}}f_{m}
-\sum_{k=N_{1}}^{N_{2}}f_{k}^{2}.
\label{A.13}
\end{equation}
Manipulating with the first term on the extreme right-hand side of
Eq.\ (\ref{A.13}), we have
\begin{equation}
\sum_{k=N_{1}}^{N_{2}}f_{k}\sum_{m=N_{1}}^{k}f_{m}
=\sum_{m=N_{1}}^{N_{2}}f_{m}\sum_{k=m}^{N_{2}}f_{k}
=\sum_{k=N_{1}}^{N_{2}}f_{k}\sum_{m=k}^{N_{2}}f_{m}.
\label{A.14}
\end{equation}
Plugging the result (\ref{A.14}) into Eq.\ (\ref{A.13}), we obtain
\begin{equation}
\left(\sum_{k=N_{1}}^{N_{2}}f_{k}\right)^{2}
=2\sum_{k=N_{1}}^{N_{2}}f_{k}\sum_{m=k}^{N_{2}}f_{m}
-\sum_{k=N_{1}}^{N_{2}}f_{k}^{2},
\label{A.15}
\end{equation}
from which the identity in Eq.\ (\ref{A.10}) follows immediately.
%
%
\subsection{The summation formula (\ref{2.46})}
\label{A.III}
\setcounter{equation}{0}
We denote
\begin{equation}
S_{3}=\sum_{k=0}^{n-1}\frac{(2k+1)^{2}}{(n-k)^{2}(n+k+1)^{2}}.
\label{A.16}
\end{equation}
If we carry out the partial fraction decomposition of the summand, we
have
\begin{equation}
S_{3}=\sum_{k=0}^{n-1}\frac{1}{(n-k)^{2}}
+\sum_{k=0}^{n-1}\frac{1}{(n+k+1)^{2}}
-\frac{2}{2n+1}\sum_{k=0}^{n-1}\frac{1}{n-k}
-\frac{2}{2n+1}\sum_{k=0}^{n-1}\frac{1}{n+k+1}
\label{A.17}
\end{equation}
and further, after obvious rearrangements,
\begin{equation}
S_{3}=\sum_{k=0}^{2n-1}\frac{1}{(k+1)^{2}}
-\frac{2}{2n+1}\sum_{k=1}^{2n}\frac{1}{k}.
\label{A.18}
\end{equation}
Now, it holds that
\begin{equation}
\sum_{k=0}^{N-1}\frac{1}{(k+1)^{2}}
=\sum_{k=0}^{\infty}\frac{1}{(k+1)^{2}}
-\sum_{k=0}^{\infty}\frac{1}{(k+N+1)^{2}}
=\psi_{1}(1)-\psi_{1}(N+1)
\qquad (N\in\mathbb{N}_{0}),
\label{A.19}
\end{equation}
where $\psi_{1}(z)=\mathrm{d}\psi(z)/\mathrm{d}z$ is the trigamma
function. On employing Eqs.\ (\ref{A.19}) and (\ref{A.5}) in Eq.\
(\ref{A.18}), after using the well-known relations
\begin{equation}
\psi(1)=-\gamma
\label{A.20}
\end{equation}
(here and below $\gamma$ stands for the Euler--Mascheroni constant)
and
\begin{equation}
\psi_{1}(1)=\frac{\pi^{2}}{6},
\label{A.21}
\end{equation}
we finally obtain
\begin{equation}
\sum_{k=0}^{n-1}\frac{(2k+1)^{2}}{(n-k)^{2}(n+k+1)^{2}}
=\frac{\pi^{2}}{6}-\frac{2\gamma}{2n+1}
-\frac{2}{2n+1}\psi(2n+1)-\psi_{1}(2n+1),
\label{A.22}
\end{equation}
which is Eq.\ (\ref{2.46}).
%
%
\subsection{The summation formula (\ref{2.47})}
\label{A.IV}
\setcounter{equation}{0}
We denote
\begin{equation}
S_{4}=\sum_{k=0}^{n-1}(-1)^{n+k}\frac{(2n+1)(2k+1)}
{(n-k)^{2}(n+k+1)^{2}}.
\label{A.23}
\end{equation}
A partial-fraction decomposition of the summand gives
\begin{equation}
S_{4}=\sum_{k=0}^{n-1}\frac{(-1)^{n+k}}{(n-k)^{2}}
-\sum_{k=0}^{n-1}\frac{(-1)^{n+k}}{(n+k+1)^{2}}.
\label{A.24}
\end{equation}
With a little bit of algebra on the right-hand side of Eq.\
(\ref{A.24}), we obtain
\begin{equation}
S_{4}=\sum_{k=0}^{2n-1}\frac{(-1)^{k+1}}{(k+1)^{2}}
\label{A.25}
\end{equation}
and further
\begin{equation}
S_{4}=-\sum_{k=0}^{2n-1}\frac{1}{(k+1)^{2}}
+\frac{1}{2}\sum_{k=0}^{n-1}\frac{1}{(k+1)^{2}}.
\label{A.26}
\end{equation}
From this, with reference to Eqs.\ (\ref{A.19}) and (\ref{A.21}),
we eventually arrive at
\begin{equation}
\sum_{k=0}^{n-1}(-1)^{n+k}\frac{(2n+1)(2k+1)}{(n-k)^{2}(n+k+1)^{2}}
=-\frac{\pi^{2}}{12}+\psi_{1}(2n+1)-\frac{1}{2}\psi_{1}(n+1),
\label{A.27}
\end{equation}
which is Eq.\ (\ref{2.47}).
%
%

%
\end{document}